OpenStudy (anonymous):

integrate1/cosx+sinx

OpenStudy (anonymous):

Do you mean \[\int\limits_{}^{} 1 / (\sin x + \cos x)\]

OpenStudy (anonymous):

\[1-2\sin ^{2}x=u , 2\cos ^{2}x-1=v\] \[du=-2(2)cosdx=-4cosdx, dv=-4sindx\] substitute in the equation.

OpenStudy (anonymous):

do you think it works??

OpenStudy (anonymous):

yes it should work.... u will get cosx=-1/4 du and sinx=-1/4dv

OpenStudy (anonymous):

so what??? @aryandecoolest ? are you kidding?

OpenStudy (anonymous):

so 1/4 (du)/u - Int (-1/4) dv/v which gives us 1/4 ln (2cos2x-1)-1/4 ln (1-2sin2x) +C

OpenStudy (anonymous):

:?????????????????????????????

OpenStudy (anonymous):

Int dx/(sinx+cosx)= Int (cosx-Sinx)dx/(Cos2x-sin2x)= Int cosx dx/ (1-2Sin2x) - Int sinx dx/ (2cos2x-1) ..(.1) Let 1-2sin2x =u and 2cos2x-1 =v so, du= -2 (2)cosx dx =-4cosx dx, thus cosx dx = -1/4 du........(2) and dv= -2 (2) sinx dx = -4sin x dx, thus sinx dx= -1/4 dv....(3) let's now plug-in the value of sin x and cosx in (1) We have int -1/4 (du)/u - Int (-1/4) dv/v =-1/4 ln(u) -1/4 ln (v)+C = -1/4 ln (2cos2x-1)-1/4 ln (1-2sin2x) +C

OpenStudy (anonymous):

This can be done with the half-angle substitution: \(t=\tan\dfrac{x}{2}\). This gives you a list of facts: \[\sin x=\sin\left(2\cdot\frac{x}{2}\right)=2\sin\frac{x}{2}\cos\frac{x}{2}\\ \cos x=\cos\left(2\cdot\frac{x}{2}\right)=\cos^2\frac{x}{2}-\sin^2\frac{x}{2}\\ dt=\frac{1}{2}\sec^2\frac{x}{2}~dx~~\iff~~2\cos^2\frac{x}{2}~dt=dx\] Given that \(t=\tan\dfrac{x}{2}\), you can set up a triangle:  so that \[\sin x=\frac{2t}{1+t^2}\\ \cos x=\frac{1-t^2}{1+t^2}\\ \frac{2}{1+t^2}~dt=dx\] We then equate the integrals under this substitution: \[\int\frac{dx}{\cos x+\sin x}={\Huge\int}\frac{\dfrac{2}{1+t^2}}{\dfrac{2t}{1+t^2}+\dfrac{1-t^2}{1+t^2}}~dt\] Simplifying gives \[\int\frac{2}{2t+1-t^2}~dt\] Completing the square in the denominator gives \[\int\frac{2}{2-(t-1)^2}~dt\] which can be taken care of with a trig sub: \(t-1=\sqrt2\sin u\), so that \(dt=\sqrt 2\cos u~du\). \[\int\frac{2\sqrt2\cos u}{2-(\sqrt2\sin u)^2}~du\\ \sqrt2\int\frac{\cos u}{1-\sin^2 u}~du\\ \sqrt2\int\frac{\cos u}{\cos^2 u}~du\\ \sqrt2\int\frac{du}{\cos u}\\ \sqrt2\int\sec u~du\]